# Multimodal foundation world models for generalist embodied agents

### v1

@@ Abstract
Learning generalist embodied agents, able to solve multitudes of tasks in different domains is a long-standing problem. Reinforcement learning (RL) is hard to scale up as it requires a complex reward design for each task. In contrast, language can specify tasks in a more natural way. Current foundation vision-language models (VLMs) generally require fine-tuning or other adaptations to be functional, due to the significant domain gap. However, the lack of multimodal data in such domains represents an obstacle toward developing foundation models for embodied applications. In this work, we overcome these problems by presenting multimodal foundation world models, able to connect and align the representation of foundation VLMs with the latent space of generative world models for RL, without any language annotations. The resulting agent learning framework, GenRL, allows one to specify tasks through vision and/or language prompts, ground them in the embodied domain’s dynamics, and learns the corresponding behaviors in imagination. As assessed through large-scale multi-task benchmarking, GenRL exhibits strong multi-task generalization performance in several locomotion and manipulation domains. Furthermore, by introducing a data-free RL strategy, it lays the groundwork for foundation model-based RL for generalist embodied agents.
Figure 1: Multimodal foundation world models connect and align the video-language space of a foundation model with the latent space of a generative world model for reinforcement learning, requiring vision-only data. Our GenRL framework turns visual and/or language prompts into latent targets and learns to realize the corresponding behaviors by training in the world model’s imagination.
Project website:mazpie.github.io/genrl
@@ 1 Introduction
Foundation models are large pre-trained models endowed with extensive knowledge of the world, which can be readily adapted for a given task [41]. These models have demonstrated extraordinary generalization capabilities in a wide range of vision [27, 42, 62] and language tasks [40, 18, 50, 11]. As we aim to extend this paradigm to embodied applications, where agents physically interact with objects and other agents in their environment, we require generalist agents that are capable of reasoning about these interactions and executing action sequences within these settings [57].
Reinforcement learning (RL) allows agents to learn complex behaviors from visual and/or proprioceptive inputs [17, 25, 26] by maximizing a specified reward function. Scaling up RL to multiple tasks and embodied environments remains challenging as designing reward functions is a complicated process, requiring expert knowledge and prone to errors which can lead to undesired behaviors [1]. Recent work has proposed the adoption of visual-language models (VLMs) to specify rewards for visual environments using language [4, 45], e.g. using the similarity score computed by CLIP [42] between an agent’s input images and text prompts. However, these approaches require fine-tuning of the VLM [37] or adaptation of the visual domain [45], to work reliably.
@@
In this work, we present GenRL, a novel approach for training generalist agents from visual or language prompts, requiring no language annotations. GenRL learns multimodal foundation world models (MFWMs), where the joint embedding space of a foundation video-language model [54] is connected and aligned with the representation of a generative world model for RL [22], using unimodal vision-only data. The MFWM allows the specification of tasks by grounding language or visual prompts into the RL domain. Then, we introduce an RL objective for learning to accomplish the specified tasks in imagination [23], by matching the prompts in latent space.
Compared to previous work in world models and VLMs for RL, one emergent property of GenRL is the possibility to generalize to new tasks in a completely data-free manner. After training the MFWM, it possesses both strong priors over the dynamics of the environment, and large-scale multimodal knowledge. This combination enables the agent to interpret a large variety of task specifications and learn the corresponding behaviors. Thus, analogously to foundation models for vision and language, GenRL allows generalization to new tasks without additional data and lays the groundwork for foundation models in embodied RL domains [41].  [^1] The code, datasets and trained models will be made publicly available.
@@ 2 Preliminaries and background
Problem setting. The agent receives from the environment observations $x\in\mathcal{X}$ and interacts with it through actions $a\in\mathcal{A}$. The objective of the agent is to accomplish a certain task $\tau$, which can be specified either in the observation space $x_{\tau}$, e.g. through images or videos, or in language space $y_{\tau}$, where $\mathcal{Y}$ represents the space of all possible sentences. Crucially, compared to a standard RL setting, we do not assume that a reward signal is available to solve the task. When a reward function exists, it is instead used to evaluate the agent’s performance.
Generative world models for RL. In model-based RL, the optimization of the agent’s actions is done efficiently, by rolling out and scoring imaginary trajectories using a (learned) model of the environment’s dynamics. In recent years, this paradigm has grown successful thanks to the adoption of generative world models, which learn latent dynamics by self-predicting the agent’s inputs [22]. World models have shown impressive performance in vision-based environments [23], improving our ability to solve complex and open-ended tasks [25]. Generative world models have been successfully extended to many applications, such as exploration [48], skill learning [39], solving long-term memory tasks [47], and robotics [55].
Foundations models for RL. Large language models (LLMs) have been used for specifying behaviors using language [38, 28, 53], but this generally assumes the availability of a textual interface with the environment or that observations and/or actions can be translated to the language domain. The adoption of vision-language models (VLMs) reduces these assumptions, as it allows the evaluation of behaviors in the visual space. However. this approach has yet to show robust performance, as it generally requires fine-tuning of the VLM [4, 15], prompt hacking techniques [9] or visual modifications to the environment [4].
Vision-language generative modelling. Given the large success of image-language generative models [46], recent efforts in the community have focused on replicating and extending such success to the video domain, where the temporal dimension introduces new challenges, such as temporal consistency and increased computational costs [29, 3]. Video generative models are similar to world models for RL, with the difference that generation models outputs are typically not conditioned on actions, but rather conditioned on language[29] or on nothing at all (i.e. an unconditional model).
@@ 3 GenRL
Figure 2: Overview of GenRL. The agent learns a multimodal foundation world model that connects and aligns (a) the representation of a foundation VLM with the latent states of a generative world model. Given a certain task prompt, (b) the model allows embedding the task and translating into targets in the latent dynamics space, which the agent can learn to achieve by using RL in imagination.
@@ 3.1 World models for RL
GenRL learns a task-agnostic world model representation by modelling the sequential dynamics of the environment in a compact discrete latent space $S$ [23, 25]. Latent states $s\in S$ are sampled from independent categorical distributions. The gradients for training the model are propagated through the sampling process with straight-through estimation [5].
@@
where $p_{\phi}(s_{t}|s_{t-1},a_{t-1})$ is a shorthand for $p_{\phi}(s_{t}|f_{\phi}(s_{t-1},a_{t-1},h_{t-1}))$. The sequence model is implemented as a linear GRU cell [8]. Differently from recurrent state space models (RSSM; [24]), for our framework, encoder and decoder models are not conditioned on the information present in the sequence model. This ensures that the latent states only contain information about a single observation. We can then leverage the encoder as a probabilistic visual tokenizer, that is grounded in the target embodied environment.
3.2 Multimodal foundation world models
@@
Vision $\begin{split}\textrm{Vision embedder:}\quad e^{(v)}=f^{(v)}_{\textrm{PT}}(x_{t:t+k}),\end{split}\quad\quad\begin{split}\textrm{Language embedder:}\quad e^{(l)}=f^{(l)}_{\textrm{PT}}(y),\end{split}$
where $x_{t:t+k}$ is a sequence of visual observations and $y$ is a text prompt. For video-language models, $k$ is generally a constant number of frames (e.g. $k\in\{4,8,16\}$ frames). Image-language models are a special case where $k=1$ as the vision embedder takes a single frame as an input. For our implementation, we adopt the InternVideo2 video-language model [54].
@@
where $\textrm{sg}(\cdot)$ indicates to stop gradients propagating.
@@
The connector learns to predict the latent states of the world model from embeddings in the VLM’s representation space. The connector’s objective consist of minimizing the KL divergence between its predictions and the world model’s encoder distribution. While more expressive architectures, such as transformers [52] or state-space models [20] could be adopted, we opt for a simpler GRU-based architecture for video modelling. This way, we keep the method simple and the architecture of the connector is symmetric with respect to the world model’s components.
@@
Multimodal VLMs trained with contrastive learning exhibit a multimodality gap [33], where the spherical embeddings of different modalities are not aligned. The role of the aligner is to reduce this multimodality gap, by projecting text embeddings into their corresponding visual embeddings. Given a dataset of vision-language data, this projective function can be learned.
In embodied domains vision-language data is typically unavailable. Instead, we leverage the idea that language embeddings can be treated as a ‘corrupted’ version of their vision counterparts [61, 63]. This allows us to approximate the language embedding with the corresponding vision embedding: $e^{(l)}\approx e^{(v)}+\epsilon$.
@@
Previous methods inject the noise into the embeddings when training the new module [61, 63], in our case the connector. However, we opt for training a separate aligner network. This allows us to train a noise-free connector, which has two main advantages: (i) it yields higher prediction accuracy for visual embedding inputs while maintaining a similar alignment for language embedding inputs; and (ii) it is more flexible; it’s easier to re-train/adapt for different noise levels, as it only requires re-training the aligner module, and its use can be avoided if unnecessary.
3.3 Learning specified task behaviors in imagination
@@
World models can be used to imagine trajectories in latent space, using the dynamics model. This allows us to train behavior policies in a model-based RL fashion [23]. Given a task specified through a visual or language prompt, our MFWM can generate the corresponding latent states by turning the embedder’s output, $e_{\textrm{task}}$, into sequences of latent states $s_{t:t+k}$ (examples are shown in Figure 1). The objective of the policy model $\pi_{\theta}$ is then to match the goals specified by the user by performing trajectory matching.
@@
The trajectory matching problem can be solved as a divergence minimization problem [13], between the distribution of the states visited by the policy $\pi_{\theta}$ and the trajectory generated using the aligner-connector networks from the user-specified prompt:
@@
The KL divergence is a natural candidate for the distance function [13]. However, in practice, we found that using the cosine distance between linear projections of the latent states notably speeds up learning and enhances stability. We can then turn the objective in Eq. 2 into a reward for RL:
@@
where $g_{\phi}$ represents the first linear layer of the world model’s decoder. We train an actor-critic model to maximize this reward and achieve the tasks specified by the user [25]. Additional implementation details are provided in Appendix B.
Temporal alignment. One issue with trajectory matching is that it assumes that the distribution of states visited by the agent starts from the same state as the target distribution. However, the initial state generated by the connector may differ from the initial state where the policy is currently in. For example, consider the Stickman agent on the right side of Figure 1. If the agent is lying on the ground and tasked to run, the number of steps to get up and reach running states may surpass the temporal span recognized by the VLM (e.g. typically 4, 8, or 16 frames), causing disalignment in the reward.
To address this initial condition alignment issue, we propose a best matching trajectory technique, inspired by best path decoding in speech recognition [19]. Our technique involves two steps:
@@
In all experiments, we fix $b=8$ (number of frames of the VLM we use [54]), which we found to strike a good compromise between comparing only the initial state ($b=1$) and performing no alignment ($b=\textrm{imagination horizon})$. An ablation study can be found in Appendix E.
@@ 4 Experiments
Overall, we employ a set of 4 locomotion environments (Walker, Cheetah, Quadruped, and a newly introduced Stickman environment) [51] and one manipulation environment (Kitchen) [21], for a total of 35 tasks where the agent is trained without rewards, only from text prompts. This represents the first large-scale study of multi-task generalization from language in RL. Details about datasets, tasks, and prompts used can be found in the Appendix C.
@@ 4.1 Offline RL
In offline RL, the objective of the agent is to learn to extract a certain task behavior from a given fixed dataset [32]. The performance of the agent generally depends on its ability to ‘retrieve’ the correct behaviors in the dataset and interpolate among them. Popular techniques for offline RL include off-policy RL methods, such as TD3 [17], advantage-weighted behavior cloning, such as IQL [30], and behavior-regularized approaches, such as CQL [31] or TD3+BC [16].
We aim to assess the multi-task capabilities of different approaches for designing rewards using VLMs. We collected large datasets for each of the domains evaluated, containing a mix of structured data (i.e. the replay buffer of agent learning to perform some tasks) and unstructured data (i.e. exploration data collected using [48]). We have removed the explicit reward information about the task and replaced it with a short task description, in language form.
@@
  • Image-language rewards: following [45], the cosine similarity between the embedding for the language prompt and the embedding for the agent’s visual observation is used as a reward. For the VLM, we adopt the SigLIP-B [60] model as it’s reported to have superior performance than the original CLIP [42].
@@
  • Video-language rewards: similar to the image-language rewards, with the difference that the vision embedding is computed from a video of the history of the last $k$ frames, as done in [15]. For the VLM, we use the InternVideo2 model [54], the same used for GenRL.
We test both approaches with a variety of offline RL methods, including IQL, TD3+BC, and TD3. All methods are trained for 500k gradient steps, and evaluated on 20 episodes. Other training details are reported in Appendix D.

### v2

# GenRL: Multimodal-foundation world models for generalization in embodied agents

@@ Abstract
Learning generalist embodied agents, able to solve multitudes of tasks in different domains is a long-standing problem. Reinforcement learning (RL) is hard to scale up as it requires a complex reward design for each task. In contrast, language can specify tasks in a more natural way. Current foundation vision-language models (VLMs) generally require fine-tuning or other adaptations to be adopted in embodied contexts, due to the significant domain gap. However, the lack of multimodal data in such domains represents an obstacle to developing foundation models for embodied applications. In this work, we overcome these problems by presenting multimodal-foundation world models, able to connect and align the representation of foundation VLMs with the latent space of generative world models for RL, without any language annotations. The resulting agent learning framework, GenRL, allows one to specify tasks through vision and/or language prompts, ground them in the embodied domain’s dynamics, and learn the corresponding behaviors in imagination. As assessed through large-scale multi-task benchmarking in locomotion and manipulation domains, GenRL enables multi-task generalization from language and visual prompts. Furthermore, by introducing a data-free policy learning strategy, our approach lays the groundwork for foundational policy learning using generative world models.
Figure 1: Multimodal-foundation world models connect and align the video-language space of a foundation model with the latent space of a generative world model for reinforcement learning, requiring vision-only data. Our GenRL framework turns visual and/or language prompts into latent targets and learns to realize the corresponding behaviors by training in the world model’s imagination.
Website, code and data:mazpie.github.io/genrl
@@ 1 Introduction
Foundation models are large pre-trained models endowed with extensive knowledge of the world, which can be readily adapted for a given task Bommasani2022FoundationModels . These models have demonstrated extraordinary generalization capabilities in a wide range of vision Kirillov2023SAM ; Radford2021CLIP ; Zhao2024VideoPrism and language tasks OpenAI2023GPT4 ; Google2024Gemini ; Touvron2023Llama ; Devlin2019BERT . As we aim to extend this paradigm to embodied applications, where agents physically interact with objects and other agents in their environment, we require generalist agents that are capable of reasoning about these interactions and executing action sequences within these settings (Yang2023FMDM, ).
Reinforcement learning (RL) allows agents to learn complex behaviors from visual and/or proprioceptive inputs Fujimoto2018TD3 ; Hafner2023DreamerV3 ; Hansen2024TDMPC2 by maximizing a specified reward function. Scaling up RL to multiple tasks and embodied environments remains challenging as designing reward functions is a complicated process, requiring expert knowledge and prone to errors which can lead to undesired behaviors Amodei2016RewardHacking . Recent work has proposed the adoption of visual-language models (VLMs) to specify rewards for visual environments using language Baumli2023VLMRew ; Rocamonde2023VLMRM ; Luo2024TextAwareDP , e.g. using the similarity score computed by CLIP Radford2021CLIP between an agent’s input images and text prompts. However, these approaches mostly require fine-tuning of the VLM Lubana2023FoMo , otherwise, they tend to work reliably only in a few visual settings Rocamonde2023VLMRM .
@@
In this work, we present GenRL, a novel approach requiring no language annotations that allows training agents to solve multiple tasks from visual or language prompts. GenRL learns multimodal-foundation world models (MFWMs), where the joint embedding space of a foundation video-language model Wang2024InternVideo2 is connected and aligned with the representation of a generative world model for RL Ha2018WM , using only vision data. The MFWM allows the specification of tasks by grounding language or visual prompts into the embodied domain’s dynamics. Then, we introduce an RL objective that enables learning to accomplish the specified tasks in imagination Hafner2020DreamerV1 , by matching the prompts in latent space.
Compared to previous work in world models and VLMs for RL, one emergent property of GenRL is the possibility to generalize to new tasks in a completely data-free manner. After training the MFWM, it possesses both strong priors over the dynamics of the environment, and large-scale multimodal knowledge. This combination enables the agent to interpret a large variety of task specifications and learn the corresponding behaviors. Thus, analogously to foundation models for vision and language, GenRL allows generalization to new tasks without additional data and lays the groundwork for foundation models in embodied RL domains Bommasani2022FoundationModels .
@@ 2 Preliminaries and background
Problem setting. The agent receives from the environment observations $x\in\mathcal{X}$ and interacts with it through actions $a\in\mathcal{A}$. In this work, we focus on visual reinforcement learning, so observations are images of the environment. The objective of the agent is to accomplish a certain task $\tau$, which can be specified either in the observation space $x_{\tau}$, e.g. through images or videos, or in language space $y_{\tau}$, where $\mathcal{Y}$ represents the space of all possible sentences. Crucially, compared to a standard RL setting, we do not assume that a reward signal is available to solve the task. When a reward function exists, it is instead used to evaluate the agent’s performance.
Generative world models for RL. In model-based RL, the optimization of the agent’s actions is done efficiently, by rolling out and scoring imaginary trajectories using a (learned) model of the environment’s dynamics. In recent years, this paradigm has grown successful thanks to the adoption of generative world models, which learn latent dynamics by self-predicting the agent’s inputs Ha2018WM . World models have shown impressive performance in vision-based environments Hafner2020DreamerV1 , improving our ability to solve complex and open-ended tasks Hafner2023DreamerV3 . Generative world models have been successfully extended to many applications, such as exploration Sekar2020Plan2Explore , skill learning Mazzaglia2023Choreographer , solving long-term memory tasks Reza2024R2I , and robotics Wu2022DayDreamer ; Ferraro2023FOCUS .
Foundations models for RL. Large language models (LLMs) have been used for specifying behaviors using language Ma2024EUREKA ; Klissarov2023Motif ; Wang2023VOYAGER ; Xie2024Text2reward , but this generally assumes the availability of a textual interface with the environment or that observations and/or actions can be translated to the language domain. The adoption of vision-language models (VLMs) reduces these assumptions, as it allows the evaluation of behaviors in the visual space. However. this approach has yet to show robust performance, as it generally requires fine-tuning of the VLM Baumli2023VLMRew ; Fan2022MineDojo , prompt hacking techniques Cui2022ZeroShot or visual modifications to the environment Baumli2023VLMRew .
Vision-language generative modelling. Given the large success of image-language generative models Rombach2022StableDiffusion , recent efforts in the community have focused on replicating and extending such success to the video domain, where the temporal dimension introduces new challenges, such as temporal consistency and increased computational costs Kondratyuk2023Videopoet ; Bartal2024Lumiere . Video generative models are similar to world models for RL, with the difference that generation models outputs are typically not conditioned on actions, but rather conditioned on language Kondratyuk2023Videopoet or on nothing at all (i.e. an unconditional model).
@@ 3 GenRL
Figure 2: Overview of GenRL. The agent learns a multimodal-foundation world model that connects and aligns (a) the representation of a foundation VLM with the latent states of a generative world model. Given a certain task prompt, (b) the model allows embedding the task and translating into targets in the latent dynamics space, which the agent can learn to achieve by using RL in imagination.
@@ 3.1 World models for RL
GenRL learns a task-agnostic world model representation by modelling the sequential dynamics of the environment in a compact discrete latent space $S$ Hafner2020DreamerV1 ; Hafner2023DreamerV3 . Latent states $s\in S$ are sampled from independent categorical distributions. The gradients for training the model are propagated through the sampling process with straight-through estimation Bengio2013STGradients .
@@
where $p_{\phi}(s_{t}|s_{t-1},a_{t-1})$ is a shorthand for $p_{\phi}(s_{t}|f_{\phi}(s_{t-1},a_{t-1},h_{t-1}))$. The sequence model is implemented as a linear GRU cell Chung2014GRU . Differently from recurrent state space models (RSSM; Hafner2019Planet ), for our framework, encoder and decoder models are not conditioned on the information present in the sequence model. This ensures that the latent states only contain information about a single observation, while temporal information is stored in the hidden state of the sequence model. Given the simpler encoder-decoder strategy of our model, the encoder can be seen as a probabilistic visual tokenizer, which is grounded in the target embodied environment Yu2024MagVit .
3.2 Multimodal-foundation world models
@@
$\begin{split}\textrm{Vision embedder:}\quad e^{(v)}=f^{(v)}_{\textrm{PT}}(x_{t:t+k}),\end{split}\quad\quad\begin{split}\textrm{Language embedder:}\quad e^{(l)}=f^{(l)}_{\textrm{PT}}(y),\end{split}$
where $x_{t:t+k}$ is a sequence of visual observations and $y$ is a text prompt. For video-language models, $k$ is generally a constant number of frames (e.g. $k\in\{4,8,16\}$ frames). Image-language models are a special case where $k=1$ as the vision embedder takes a single frame as an input. For our implementation, we adopt the InternVideo2 video-language model Wang2024InternVideo2 (with $k$=8).
@@
[…]
@@
The connector learns to predict the latent states of the world model from embeddings in the VLM’s representation space. The connector’s objective consists of minimizing the KL divergence between its predictions and the world model’s encoder distribution. While more expressive architectures, such as transformers Vaswani2023Transformers or state-space models Gu2023Mamba could be adopted, we opt for a simpler GRU-based architecture for video modelling. This way, we keep the method simple and the architecture of the connector is symmetric with respect to the world model’s components.
Aligning multimodal representations. The connector learns to map visual embeddings from the pretrained VLM to latent states of the world model. When learning the connector from visual embeddings $e^{(v)}$, we assume it can generalize to the (theoretical) corresponding language embedding $e^{(l)}$ if the angle $\theta$ between the two embeddings is small enough, as shown in Fig. 3a. This can be expressed as $\cos{\theta}>c$ or $\theta<\arccos{c}$, with $c$ a small positive constant Zhou2022LAFITE .
@@
Multimodal VLMs trained with contrastive learning exhibit a multimodality gap Liang2022MMGap , where the spherical embeddings of different modalities are not aligned. Given a dataset of vision-language data, this projective function can be learned. However, in embodied domains vision-language data is typically unavailable. Thus, we have to find a way to align the representations using no language annotations.
Previous methods inject the noise into the vision embeddings during training Zhang2024C3 ; Zhou2022LAFITE . This leads to the situation shown in Figure 3b, where $c$ grows larger with the noise. This allows language embeddings to be close enough to their visual counterparts.
In our work, we instead learn an aligner network, which maps points surrounding $e^{(v)}$ closer to $e^{(v)}$. As represented in Fig. 3c, this way, $c$ is unaltered but the aligner will map $e^{(l)}$ close enough to $e^{(v)}$. Since we use noise to sample points around $e^{(v)}$ the aligner model can be trained using vision-only data and thus, no language annotations.
Figure 3: When training the connector on (a) the VLM’s representation we can address the multimodality gap in multiple ways: (b) prior works adopt noise during the training of the connector, (c) we adopt an aligner network that learns to map points in proximity of the visual embedding close the corresponding embedding.
@@
The aligner allows us to train a noise-free connector, which has two main advantages: (i) it yields higher prediction accuracy for visual embedding inputs while maintaining a similar alignment for language embedding inputs; and (ii) it is more flexible; it’s easier to re-train/adapt for different noise levels, as it only requires re-training the aligner module, and its use can be avoided if unnecessary.
3.3 Specifying and learning tasks in imagination
@@
World models can be used to imagine trajectories in latent space, using the sequential and dynamics models. This allows us to train behavior policies in a model-based RL fashion Hafner2020DreamerV1 . Given a task specified through a visual or language prompt, our MFWM can generate the corresponding latent states by turning the embedder’s output, $e_{\textrm{task}}$, into sequences of latent states $s_{t:t+k}$ (decoded examples are shown in Figure 1). The objective of the policy model $\pi_{\theta}$ is then to match the goals specified by the user by performing trajectory matching.
@@
The trajectory matching problem can be solved as a divergence minimization problem Englert2013MBTrajectoryMatching , between the distribution of the states visited by the policy $\pi_{\theta}$ and the trajectory generated using the aligner-connector networks from the user-specified prompt:
@@
The KL divergence is a natural candidate for the distance function Englert2013MBTrajectoryMatching . However, in practice, we found that using the cosine distance between linear projections of the latent states notably speeds up learning and enhances stability. We can then turn the objective in Eq. 2 into a reward for RL:
@@
where $g_{\phi}$ represents the first linear layer of the world model’s decoder. We train an actor-critic model to maximize this reward and achieve the tasks specified by the user Hafner2023DreamerV3 . Additional implementation details are provided in Appendix B.
Temporal alignment. One issue with trajectory matching is that it assumes that the distribution of states visited by the agent starts from the same state as the target distribution. However, the initial state generated by the connector may differ from the initial state where the policy is currently in. For example, consider the Stickman agent on the right side of Figure 1. If the agent is lying on the ground and tasked to run, the number of steps to get up and reach running states may surpass the temporal span recognized by the VLM (e.g. in our case 8 frames), causing disalignment in the reward.
To address this initial condition alignment issue, we propose a best matching trajectory technique, inspired by best path decoding in speech recognition Graves2006CTC . Our technique involves two steps:
@@
In all experiments, we fix $b=8$ (number of frames of the VLM we use Wang2024InternVideo2 ), which we found to strike a good compromise between comparing only the initial state ($b=1$) and performing no alignment ($b=\textrm{imagination horizon})$. An ablation study can be found in Appendix E.
@@ 4 Experiments
Overall, we employ a set of 4 locomotion environments (Walker, Cheetah, Quadruped, and a newly introduced Stickman environment) tunyasuvunakool2020dmcontrol and one manipulation environment (Kitchen) gupta2019kitchen , for a total of 35 tasks where the agent is trained without rewards, using only visual or language prompts. Details about the datasets, tasks, and prompts used can be found in the Appendix C.
@@ 4.1 Offline RL
In offline RL, the objective of the agent is to learn to extract a certain task behavior from a given fixed dataset Levine2020OfflineRL . The performance of the agent generally depends on its ability to ‘retrieve’ the correct behaviors in the dataset and interpolate among them. Popular techniques for offline RL include off-policy RL methods, such as TD3 Fujimoto2018TD3 , advantage-weighted behavior cloning, such as IQL Kostrikov2021IQL , and behavior-regularized approaches, such as CQL Kumar2020CQL or TD3+BC Fujimoto2021TD3+BC .
We aim to assess the multi-task capabilities of different approaches for designing rewards using VLMs. We collected large datasets for each of the domains evaluated, containing a mix of structured data (i.e. the replay buffer of an agent Hafner2023DreamerV3 learning to perform some tasks) and unstructured data (i.e. exploration data collected using Sekar2020Plan2Explore ). The datasets contain no reward information and no text annotations of the trajectories. The rewards for training for a given task must be inferred by the agent, i.e. using the cosine similarity between observations and the given prompt or, in the case of GenRL, using our reward formulation (Eq. 3).
@@
  • Image-language rewards: following Rocamonde2023VLMRM , the cosine similarity between the embedding for the language prompt and the embedding for the agent’s visual observation is used as a reward. For the VLM, we adopt the SigLIP-B Zhai2023SigLIP model as it’s reported to have superior performance than the original CLIP Radford2021CLIP .
@@
  • Video-language rewards: similar to the image-language rewards, with the difference that the vision embedding is computed from a video of the history of the last $k$ frames, as done in Fan2022MineDojo . For the VLM, we use the InternVideo2 model Wang2024InternVideo2 , the same used for GenRL.
The evaluation compares GenRL to various offline RL methods from the literature, including IQL, TD3+BC, and TD3. We also introduce a model-based baseline, WM-CLIP. This baseline is the antithesis of GenRL as, rather than learning a connector and an aligner, it learns a “reversed connector". This module learns to predict VLM embeddings from the world model states (GenRL does the opposite). This makes it possible to compute rewards in imagination in a similar way to the model-free baselines, by computing the cosine similarity between the visual embeddings predicted from imagined states and the task’s language embeddings.
All methods are trained for 500k gradient steps, and evaluated on 20 episodes. For each task, model-free agents require training the agent from scratch, including the visual encoder, actor, and critic networks on the entire dataset. Model-based agents require training the model once for each domain and then training an actor-critic for each task. Other training and baseline details are reported in Appendix D.